OpenStudy (montanaok):

Write the specification 14.35 ≤ d ≤ 15.15 as an absolute value inequality. A. |d – 14.75| ≥ 0.4 B. |d – 0.4| ≥ 14.75 C. |d – 14.75| ≤ 0.4 D. |d – 0.4| ≤ 14.75

OpenStudy (socccerkid21):

its C im 95.73864737% sure

563blackghost (563blackghost):

We need to first look on how far apart 14.35 and 15.15 are to do this we would subtract them from each other... \(\Large{15.15-14.35=0.8}\) Since they are only \(\large{.8}\) units from each other we would divide that by 2 and see that it would equal \(\large{.4}\) So we would want each side to be equal to \(\large{.4}\) in order to accomplish that we need to find a number that will make each side be equal of \(\large{.4}\) in this case we would use \(\large{14.75}\) so we would do this... \(\LARGE {14.35-14.75 \le x - 14.75 \le 15.15 -14.75}\) In which we simplify... \(\huge\color{green}{-0.4 \le x - 14.75 \le 0.4}\) Now we see that we have a negavtive the the left and a postive to the right we can convert this to.... \(\huge\color{red}{|x-14.75| \le 0.4}\)

563blackghost (563blackghost):

@montanaok